20220708_DX_Day DX AdSec Flat Measurement, DX LBTI Closed Loop Test

Preparation

Flats

  • Current flat
[AOeng@dxadsec Applicable]$ ls -l default.sav 
lrwxrwxrwx. 1 AOeng aoeng 68 May 13 07:24 default.sav -> /local/aomeas/adsec_calib/flat/20220422_191142/flattening_result.sav
lrwxrwxrwx. 1 AOeng aoeng 11 Sep 19  2019 LBTI.sav -> default.sav
lrwxrwxrwx. 1 AOeng aoeng 11 Mar 31  2018 LUCIFER.sav -> default.sav

20220422

10:30 Full mask is 20220419_070400.mask

11:23 Full mask (above name)
  • modes 101-200 20220422_184736
  • AdOpt > print, opt_int_mat_reduce2(200, track='20220422_180622')
  • 2022040422_191142/flattening_result.sav
    • 0.527 w 0.0933 rms (-0.14 0.01) act 217, 216 @ -0.12

20220420

09:48 aligned the M3 and hexapod to RR
  • M3:Tip:-129.016, Tilt:1003.951, Piston: 0.00, RZ:56.410
  • Hexpod:X:-3.2887, Y:2.5043, Z:-1.6669, RX:-235.5, RY:-260.3, RZ:0.00

Final m2, m3 alignment

  • m2.jpg:
    m2.jpg
  • m3.jpg:
    m3.jpg

Actuator list

  • Diff of bad_actuator.list.night/day
[AOeng@dxadsec TS3]$ diff bad_actuators.list.night bad_actuators.list.day 
67c67
<      4 :  1  :  0  :  0  : 20180118 :  DLM GB : Detached magnet on TS3 - removed from list since magnet was re-attached 08/07/19; back 20220418,XZ,GB,DM
---
> #     4 :  1  :  0  :  0  : 20180118 :  DLM GB : Detached magnet on TS3 - removed from list since magnet was re-attached 08/07/19; back 20220418,XZ,GB,DM
91c91
<   260 :  0  :  1  :  0  : 20180121 :  GB : Incorrect position, off with respect to neighbouring actuators 08/06/19 temp. back, removed 08/14/19  #actuator fixed GB/XZ Jan 24,2022;back 20220418,XZ,GB,DM
---
> #  260 :  0  :  1  :  0  : 20180121 :  GB : Incorrect position, off with respect to neighbouring actuators 08/06/19 temp. back, removed 08/14/19  #actuator fixed GB/XZ Jan 24,2022;back 20220418,XZ,GB,DM
95c95
<   218 :  0  :  1  :  0  : 20181117 :  DLM : Removed during LN dattime test 08/06/19 temp. back, removed 08/14/19; back 20220418,XZ,GB,DM
---
> #  218 :  0  :  1  :  0  : 20181117 :  DLM : Removed during LN dattime test 08/06/19 temp. back, removed 08/14/19; back 20220418,XZ,GB,DM
101c101
<   367 :  0  :  1  :  0  : 20190825 :  DLM : Drifting while flattening with 4D on time scales of ~30 minutes #GB/XZ no avaiable fix, Mar 3, 2022; back 20220418,XZ,GB,DM
---
> #  367 :  0  :  1  :  0  : 20190825 :  DLM : Drifting while flattening with 4D on time scales of ~30 minutes #GB/XZ no avaiable fix, Mar 3, 2022; back 20220418,XZ,GB,DM
111,112c111,112
<   340 :  0  :  1  :  0  : 20190918 :  DLM : jumping causing low mode flats to be applied  #GB/XZ no avaiable fix, Mar 3, 2022; back 20220418,XZ,GB,DM
<   470 :  0  :  1  :  0  : 20190918 :  DLM : jumping causing low mode flats to be applied  #GB/XZ no avaiable fix, Mar 3, 2022; back 20220418,XZ,GB,DM
---
> #  340 :  0  :  1  :  0  : 20190918 :  DLM : jumping causing low mode flats to be applied  #GB/XZ no avaiable fix, Mar 3, 2022; back 20220418,XZ,GB,DM
> #  470 :  0  :  1  :  0  : 20190918 :  DLM : jumping causing low mode flats to be applied  #GB/XZ no avaiable fix, Mar 3, 2022; back 20220418,XZ,GB,DM
120,121c120,121
<   565 :  0  :  1  :  0  : 20220308 :  GB,XZ : MMC actuator; back 20220418,XZ,GB,DM
<   555 :  0  :  1  :  0  : 20220422 :  GB,XZ,DM : given high force after slew to 30 De
---
>   555 :  0  :  1  :  0  : 20220422 :  GB,XZ,DM : given high force after slew to 30 Deg El
> #  565 :  0  :  1  :  0  : 20220308 :  GB,XZ : MMC actuator; back 20220418,XZ,GB,DM

  • Diff of bad_actuator.list/.day
[AOeng@dxadsec TS3]$ diff bad_actuators.list bad_actuators.list.day 
101c101
<   367 :  0  :  1  :  0  : 20190825 :  DLM : Drifting while flattening with 4D on time scales of ~30 minutes #GB/XZ no avaiable fix, Mar 3, 2022; MMC actuator 20220510,XZ,GB
---
> #  367 :  0  :  1  :  0  : 20190825 :  DLM : Drifting while flattening with 4D on time scales of ~30 minutes #GB/XZ no avaiable fix, Mar 3, 2022; back 20220418,XZ,GB,DM
111c111
<   340 :  0  :  1  :  0  : 20190918 :  DLM : jumping causing low mode flats to be applied  #GB/XZ no avaiable fix, Mar 3, 2022; MMC actuator 20220510,XZ,GB
---
> #  340 :  0  :  1  :  0  : 20190918 :  DLM : jumping causing low mode flats to be applied  #GB/XZ no avaiable fix, Mar 3, 2022; back 20220418,XZ,GB,DM
121,127c121,122
<   115 :  0  :  1  :  0  : 20220510 :  GB,XZ : MMC actuator
<   484 :  0  :  1  :  0  : 20220510 :  GB,XZ : MMC actuator
<   628 :  0  :  1  :  0  : 20220510 :  GB,XZ : MMC actuator
<   126 :  0  :  1  :  0  : 20220510 :  GB,XZ : MMC actuator
<   471 :  0  :  1  :  0  : 20220622 :  GB,XZ : MMC actuator
<   479 :  0  :  1  :  0  : 20220622 :  GB,XZ : MMC actuator
<   454 :  0  :  1  :  0  : 20220622 :  GB,XZ : MMC actuator
---
> #  565 :  0  :  1  :  0  : 20220308 :  GB,XZ : MMC actuator; back 20220418,XZ,GB,DM
> # 471 :  0  :  1  :  0  : 20220308 :  GB,XZ : MMC actuator
129,130c124
<   321 :  0  :  1  :  0  : 20220622 :  GB,XZ : MMC actuator
<   595 :  0  :  1  :  0  : 20220622 :  GB,XZ : MMC actuator
---
> # 321 :  0  :  1  :  0  : 20220308 :  GB,XZ : MMC actuator
134,138c128
<   565 :  0  :  1  :  0  : 20200202 :  GB,XZ :High forces limit the flat (300 modes) ##GB/XZ ffb candidate Jan 25,2022;  MMC actuator; 20220510,XZ,GB
<   239 :  0  :  1  :  0  : 20200527 :  GB : actuator at 139 um current saturated
<   357 :  0  :  1  :  0  : 20220617 :  GB/XZ : actuator found at 38um with shell rested
<   23  :  0  :  1  :  0  : 20220622 :  GB/XZ : drift after set flat
<   469 :  0  :  1  :  0  : 20220622 :  GB/XZ : drift after set flat and put in MMC
---
> #  565 :  0  :  1  :  0  : 20200202 :  GB,XZ :High forces limit the flat (300 modes) ##GB/XZ ffb candidate Jan 25,2022

  • list of actuators NOT in use for the default flat (/local/aomeas/adsec_calib/flat/20220422_191142/flattening_result.sav) - for a total of 39 actuators
    • 43 45 64 65 70 71 72 79 91 116 118 129 159 168 169 170 171 208 257 299 355 372 376 377 378 379 380 381 446 460 465 524 556 557 558 559 573 578 591
  • list of actuators NOT in the m2c (KL_v29), same as above plus 555 - - for a total of 40 actuators
  • list of actuators currently in the MMC (ordered according to their drift in um as measured during the night of 07/06/2022)
    • 367 (2.8), 340 (1.5), 469 (0.85), 479 (0.6), 565 (0.55), 471 (0.35), 454 (0.25), 484 (0.22), 115 (0.13), 595 (0.12), 628 (0.11), 126 (0.09), 321 (0.09)

NONE of which are NOT in the m2c, which means that they can all in principle be used for a new flat

  • list of actuators added in the act_wo_curr since the default flat - NONE
  • list of actuators added in the act_wo_pos since the default flat, excluding the MMC ones:
    • 23 (20220622 : GB/XZ : drift after set flat)
    • 239 (2020(2?)0527 : GB : actuator at 139 um current saturated)
    • 357 (20220617 : GB/XZ : actuator found at 38um with shell rested)
    • 555 (20220422 : GB,XZ,DM : given high force after slew to 30 Deg El) also not in the m2c

these last four should not be used for a new flat

  • DX ASM MMC actuators 20220708.jpg:
    DX ASM MMC actuators 20220708.jpg

Starting alignment of 4D

Freom 20220417

08:38 Start up 4D server software
  • $ InterferometerCtrl -i interferometerctrl
  • AdSec IDL terminal
    • AdOpt >set_plot, 'X'
    • AdOpt >plot, indgen(4)
      • Should see plot
    • AdOpt > print, idl_4d_init()
    • AdOpt > a=get_img4d(/show)
    • AdOpt >help, a
      • A FLOAT = Array[892, 892]
  • If the 4D disk is not mounted on the AdSec computer then we get the error
AdOpt>a=get_img4d()
% Variable is undefined: MASK.
% Error occurred at: GET_IMG4D          59
   /home/aoeng/soul/mixed_mode_control/aoroot/idl/adsec_lib/adsec_utilities/g
  et_img4d.pro
%                    $MAIN$          
% Execution halted at: $MAIN$          

Acquiring RAW frames from 4D interferometer
img4d
img4d
4D Capturing and Processing completed
reading file

08:47 Measure Hexapod Reconstructor
  • $ idl
  • IDL> @startup-onlycom
  • IDL> print, idl_4d_init()
  • IDL> print, hexalign(/show)
  • IDL> print, hexalign(/show,/apply)

Measurement of original flat on 20220422

11:23 Full mask (above name)
  • modes 101-200 20220422_184736
  • AdOpt > print, opt_int_mat_reduce2(200, track='20220422_18')
  • 2022040422_191142/flattening_result.sav * 0.527 w 0.0933 rms (-0.14 0.01) act 217, 216 @ -0.12

Measurement of the flat from 20220422 on 20220708

  • 1.09 PVr 0.167 rms (@632.8)
Robust Peak-Valley (PVr), originally proposed by Dr. Chris Evans of the University of North Carolina Charlotte (and formerly of NIST), provides a more robust measure of surface shape. It is calculated from the peak-to-valley of a 35-term Zernike fit to the measured data plus 3*RMS of the residual (the residual is the dataset minus the 35-term fit). The result will again be displayed in the Dataset Statistics table found throughout 4Sight.

  • flat_orig_scale_better.jpg:
    flat_orig_scale_better.jpg
  • flat_orig_opd_8_removed.jpg:
    flat_orig_opd_8_removed.jpg
  • flat_orig_better.jpg:
    flat_orig_better.jpg
  • flat_orig_better_force.jpg:
    flat_orig_better_force.jpg

17:03 Rest and Set the shell and look at actuator 148 position
  • 628 modes applied, as before
  • Force on 148 is the same 0.07N
  • 4D image looks the same as before, ~4 fringes

From 20220419

10:28 Apply a gaussian bump at actuator 148 to try to get edge up
  • AdOpt >v=gauss_comm(148, 1.0)
  • AdOpt >v[adsec.act_wo_cl] = 0
  • AdOpt >v2 = 20e-9*v (+ will move force in the negative direction)
  • AdOpt >print, fsm_apply_cmd(v2, /delta, /passed)

17:50 Find that bumping 149 and 150 in the (-) direction to remove bump at 11:00 direction
  • save flat on AdSec Control GUI flat_check_act_150_bump.sav
  • Collected and processed 5 4D images (.raw, 4D) with low 8 Zernikes removed
    • /local/aomeas/adsec_calib/4d/flat_check_act_150

18:05 Rest and Set
  • Collected and processed 5 4D images (.raw, 4D) with low 8 Zernikes removed
    • /local/aomeas/adsec_calib/4d/flat_check_orig

19:52. Mike resolved the M3 DX leg 4 (selector) following error. Sertting up for closed loop test
  • M2 Global -6.6, 2.2, -1.1, -139, -645, 0, M3 Global -1135, -80, 0, 0. Baysides -50, -37, -7
Preset AO through WFS sent with 9.5 which was too faint and caused OCAM illumination error. OCAM not live and did not return with just changing binnings. cycling the BCU2K which resolved. OCAM live again.

20:03 Resending presetAO as 6. Using the 1" disturbance with the kl29
  • Mirror looks happy but pupils look eliptical and the center of the pupil looks a little offset from the cross hairs
  • The theoretical pupil is 40.0 and measured is ~37
  • TN20220708_201122
20:27 Preset AO without disturbance file
  • 20220708_203050, 2030111, 203143 with gains of 0.15,0.05,0.05 there are squigglies across the pupils
  • 204359 adjusted cub rotator to 202.402 where lines iand vignetting in pupils were minimized but Diff Y is 0.6
  • Made new pupils 20220708_205005
  • Some vignetting around 1:30 but moving the rotator to 202.308 corrected
  • These were all with KL28. We want KL29
21:09. Preset AO with kl29 rec:20220607_190028 using pure integrator SOUL filtering
  • OCAM not live during center pupils. Rebinning resolved
  • Pupils in table_LBTI_ACE-AO.900Hz_NOBin3_20220606.txt railed in Y so using ones I just made
  • TT and HO1 with 20220607_190028 , applying Ho2 causes bad thing
  • Using the 2022060619
21:59 Preset with KL28. Optimal pupils and cube rotator at 202.308
  • 0.2, 0.1, 0.1 221733, 221829
  • petal pattern on the 222455, 222512, 222525

    • Pupils with ADC's misaligned:
      ADCs misaligned.png

    • Slopes with misaligned ADC's:
      ADCs_Misaligned_Slopes.png

22:28 Closing loop with no disturbance fixing ADC
  • adjusted the cube rotator a little to see if that helped with odd illuminization of the pupils. did not. Then rehomed the adc's. That did resolve it. Less flower petal structure now wiht closed loop.
  • tt 0.2, ho1 0.1, ho2 0.1 224017, 224047, 224151
  • Optimized 224547, 224607, 224623
    • KL_v28 closed with no disturbance. Little structure on the left:
      KL28_closed_no_disturbance.png

22:53 1" distrubance now with alligned ADC
  • tt 0.2, ho1 0.225409, 225431, 225444
  • Optimized 225654, 225714, 225728
    • KL_v28 closed with 1" disturbance:
      KL28_closed_with_disturbance.png
Conclusion: Fixing the alignment of the cube rotator and ADC's seemed to help with KL28 reduction of aberrations seen on pupils. No issues closing the loop. KL29 any reconstructors has issues when applying ho2.

-- DougMiller - 08 Jul 2022
I Attachment Action Size Date Who Comment
ADCs misaligned.pngpng ADCs misaligned.png manage 35 K 08 Jul 2022 - 23:18 JenniferPower Pupils with ADC's misaligned
DX ASM MMC actuators 20220708.jpgjpg DX ASM MMC actuators 20220708.jpg manage 318 K 08 Jul 2022 - 16:47 DougMiller  
KL28_closed_no_disturbance.pngpng KL28_closed_no_disturbance.png manage 14 K 08 Jul 2022 - 23:21 JenniferPower KL_v28 closed with no disturbance. Little structure on the left
KL28_closed_with_disturbance.pngpng KL28_closed_with_disturbance.png manage 13 K 08 Jul 2022 - 23:26 JenniferPower KL_v28 closed with 1" disturbance
flat_orig_better.jpgjpg flat_orig_better.jpg manage 144 K 08 Jul 2022 - 17:01 DougMiller  
flat_orig_better_force.jpgjpg flat_orig_better_force.jpg manage 98 K 08 Jul 2022 - 17:02 DougMiller  
flat_orig_opd_8_removed.jpgjpg flat_orig_opd_8_removed.jpg manage 90 K 08 Jul 2022 - 16:42 DougMiller  
flat_orig_scale_better.jpgjpg flat_orig_scale_better.jpg manage 171 K 08 Jul 2022 - 16:27 DougMiller  
m2.jpgjpg m2.jpg manage 17 K 08 Jul 2022 - 17:27 DougMiller  
m3.jpgjpg m3.jpg manage 21 K 08 Jul 2022 - 18:18 DougMiller  
Topic revision: r5 - 08 Jul 2022, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback